# Simulating Quantum Circuits with Tree Tensor Networks using Density-Matrix Renormalization Group Algorithm

### v3

@@ II Theory
Figure 2: DMRG algorithm for simulating quantum circuits is illustrated. (a) A schematic of the compression step: $k$ gates are applied to the initial state |$\ket{\Psi^{i}}$, followed by an approximation of the resulting state with a tree tensor network |$\ket{\widetilde{\Psi}^{i+k}}$. (b) Tensor network diagram of the scalar product being maximized; the purple node indicates the tensor currently being optimized, which serves as the orthogonality center. (c–d) The optimal tensor $\widetilde{T}^{[v]}_{\text{opt}}$ is computed from the environment tensor $F^{[v]}$, normalized by $\sqrt{\text{Tr}F^{[v]}F^{[v]*}}$. (e) The orthogonality center is shifted to the next tensor to be optimized, $\widetilde{T}^{[v^{\prime}]}$, via a sequence of QR decompositions.
@@ II.1 Tree tensor network
An arbitrary quantum state |$\ket{\Psi}$ describing a system of $N$ qubits is expressed as
|$\ket{\Psi}=\sum_{q_{1}q_{2}\dots q_{N}}\Psi_{q_{1}q_{2}\dots q_{N}}\ket{q_{1}q_{2}\dots q_{N}},$ (1)
where $\Psi_{q_{1}q_{2}\dots q_{N}}$ is the coefficient tensor and |$\ket{q_{1}q_{2}\dots q_{N}}$ denotes the basis states. The coefficient tensor requires $2^{N}$ parameters to be specified exhibiting exponential scaling with the system size. A TTN decomposes this large tensor into a product of smaller tensors with a hierarchical tree structure where the individual tensors are located at the nodes of a connected acyclic graph and the edges between nodes indicate how these tensors are contracted as shown in Fig. 1. The TTN ansatz is given as
@@ II.2 Quantum circuit simulation
To formalize our algorithm, let us consider a TTN at step $i$ denoted by |$\ket{\Psi^{i}}$, which ideally approximates the exact quantum state after the application of $i$ gates. When a sequence of $k$ additional gates, $\hat{U}_{i+k}\cdots\hat{U}_{i+1}$, is applied, the updated state is given as
|$\ket{\Psi^{i+k}}=\hat{U}_{i+k}\cdots\hat{U}_{i+1}\ket{\Psi^{i}}.$ (3)
The goal is to approximate this evolved state with a new TTN |$\ket{\widetilde{\Psi}^{i+k}}$, whose structure can be chosen arbitrarily and is independent of the previous TTN |$\ket{\Psi^{i}}$. For all the simulations in this work, we keep the structure of TTN fixed. The approximation is obtained by minimizing the distance
$\delta=\left\|\ket{\Psi^{i+k}}-\ket{\widetilde{\Psi}^{i+k}}\right\|^{2},$ (4)
through iterative optimization of each tensor $\widetilde{T}^{[v]}$ in |$\ket{\widetilde{\Psi}^{i+k}}$ individually while keeping the others fixed. Each such optimization step reduces to solving a linear system of equations, whose solution yields the optimal tensor $\widetilde{T}^{[v]}_{\text{opt}}$. To simplify this optimization, it is advantageous to put |$\ket{\widetilde{\Psi}^{i+k}}$ into a canonical form with respect to the target tensor $\widetilde{T}^{[v]}$, as illustrated in Fig. 2(b). In this form, all other tensors in the network are isometric in the direction of $\widetilde{T}^{[v]}$, which serves as the orthogonality center. This ensures that the norm of the TTN is simply given by
$\braket{\widetilde{\Psi}^{i+k}}{\widetilde{\Psi}^{i+k}}=\text{Tr}\widetilde{T}^{[v]}\widetilde{T}^{[v]*},$ (5)
where the Tr denotes the full contraction of the network with summation over all indices. The optimal target tensor (up to normalization) is then obtained from the environment tensor $F^{[v]}$, which is computed by contracting the overlap network $\braket{\widetilde{\Psi}^{i+k}}{\Psi^{i+k}}$ without the tensor $\widetilde{T}^{[v]*}$ [9], as shown in Fig 2(c–d). The resulting expression for the optimal target tensor is given as
@@
$f_{v}=\left\|\braket{\widetilde{\Psi}^{i+k}}{\hat{U}_{i+k}\cdots\hat{U}_{i+1}}{\Psi^{i}}\right\|^{2},$ (7)
@@
[…]
@@ III Results
In the following, the results of simulations for random circuits and QAOA circuits across various qubit connectivity configurations are presented. For all the simulations, the initial state is set to |$\ket{00\cdots 00}$.
@@ III.1.2 Tree-like circuit connectivity
[…]
@@ III.2 QAOA
[…]
@@
[…]
@@ III.2.2 Tree-like circuits
[…]
@@ Appendix A Comparison with SVD-based approach
In this section, we compare the DMRG method used in our work with the SVD-based truncation scheme described in Ref. [30]. The SVD-based approach approximates the quantum state by applying an SVD after each gate operation and discarding singular values below a certain threshold. These truncations introduce an error with each gate. These errors compound with subsequent gate applications, potentially leading to a significant loss of fidelity.
@@
Figure 16: Comparison of error rates $\epsilon$ as a function of bond dimension $\chi$ obtained from the simulation of random nearest-neighbor circuits using MPS and TTN. The results are for a system of $N=32$ qubits with $N_{2g}=310$ two-qubit gates. The inset shows error rates $\epsilon$ as a memory footprint $M$ obtained from the same simulation.
@@
[…]
Fig. 15 shows the same comparison, now incorporating a “naive” qubit reordering strategy as Sec. III.2. In both cases, a binary TTN is used, but the strongly correlated qubits are placed together in the TTN. While both approaches benefit from reordering, our DMRG-based algorithm at $\chi=16$ still outperforms Ref. [30] at $\chi=64$. Notably, the performance improvement is more pronounced for Ref. [30]. This improvement indicates that the SVD-based method is more sensitive to TTN structure and less robust in capturing the effect of long-range gates. This further highlights the advantage of our variational framework.
@@ Appendix B Comparison between MPS and TTN
We compare MPS and TTN as variational ansatz within the DMRG algorithm for simulating the quantum circuits. All simulations are performed on a system of $N=32$ qubits. For TTN, we adopt a binary tree structure. It is important to note that a direct comparison between TTN and MPS is inherently challenging. As a result, our analysis primarily focuses on the bond dimension $\chi$, which determines the maximum half-chain entanglement entropy that each ansatz can represent [9]. In some instances, we also provide a comparison of the memory footprint.
@@
[…]

### v4

@@ II Theory
[…]
@@ II.1 Tree tensor network
[…]
@@ II.2 Quantum circuit simulation
[…]
$\braket{\widetilde{\Psi}^{i+k}|\widetilde{\Psi}^{i+k}}=\text{Tr}\widetilde{T}^{[v]}\widetilde{T}^{[v]*},$ (5)
where the Tr denotes the full contraction of the network with summation over all indices. The optimal target tensor (up to normalization) is then obtained from the environment tensor $F^{[v]}$, which is computed by contracting the overlap network $\braket{\widetilde{\Psi}^{i+k}|\Psi^{i+k}}$ without the tensor $\widetilde{T}^{[v]*}$ [9], as shown in Fig 2(c–d). The resulting expression for the optimal target tensor is given as
@@
$f_{v}=\left\|\braket{\widetilde{\Psi}^{i+k}|\hat{U}_{i+k}\cdots\hat{U}_{i+1}|\Psi^{i}}\right\|^{2},$ (7)
@@
As demonstrated in Refs. [6, 27], $\mathcal$$\tilde${F}} serves as an accurate approximation of the exact fidelity between the exact and compressed states. Since we employ the same algorithm as in Ref. [27], this approximation remains valid in our case. Based on this, we define the error rate per two-qubit gate in our simulation as
@@ III Results
[…]
@@ III.1.2 Tree-like circuit connectivity
Figure 7: Error per gate $\epsilon$ as a function of bond dimension $\chi$ obtained from the simulation of random tree-like circuits. The results are shown for $N_{2g}=156,326,$ and $480$ two-qubit gates, corresponding to systems of $N=27,54,$ and $81$ qubits. The inset shows fidelity $\mathcal$$\tilde${F}} as a function of bond dimension $\chi$ obtained from the same simulations.
In Fig. 7, the error per gate is shown as a function of the bond dimension $\chi$ for the simulation of random tree-like circuits with system sizes of $N=27$, $54$, and $81$ qubits. The corresponding total number of two-qubit gates is $N_{2g}=156$, $316$, and $480$, respectively. Notably, an error of $\epsilon\approx 1.5\%$ is achieved even for a small bond dimension of $\chi=15$ for all the simulations. The inset displays the fidelities $\mathcal$$\tilde${F}} from the same simulations. This suggests that having similar error values $\epsilon$ across circuits with different qubit counts does not ensure their fidelities will be comparable. Therefore, as the qubit count increases, low fidelities are likely at such small bond dimensions due to the high entanglement generated in these circuits.
@@ III.2 QAOA
Figure 10: Comparison of fidelities $\mathcal$$\tilde${F}} as a function of bond dimension $\chi$ obtained from the simulation of QAOA circuits on random 3-regular graphs using TTN constructed via blind and naive approaches. The results are shown for $N=64$ at depth $p=1$ with $N_{2g}=96$ two-qubit gates.
@@
Figure 11: Fidelity $\mathcal$$\tilde${F}} as a function of bond dimension $\chi$ obtained from the simulation of QAOA circuits on bridged 3-regular graphs at circuit depth $p=4$. The results are shown for $N_{2g}=384,768,$ and $1536$ two-qubit gates, corresponding to systems of $N=64,128,$ and $256$ qubits, respectively.
@@ III.2.2 Tree-like circuits
Figure 12: Fidelity $\mathcal$$\tilde${F}} as a function of bond dimension $\chi$ obtained from the simulation of QAOA tree-like circuits at circuit depth $p=4$. The results are shown for $N_{2g}=156,316,$ and $480$ two-qubit gates, corresponding to systems of $N=27,54,$ and $81$ qubits, respectively.
@@ Appendix A Comparison with SVD-based approach
In this section, we compare the DMRG method used in our work with the SVD-based truncation scheme described in Ref. [30]. The SVD-based approach approximates the quantum state by applying an SVD after each gate operation and discarding singular values below a certain threshold. The truncations introduce an error with each gate. These errors compound with subsequent gate applications, potentially leading to a significant loss of fidelity.
@@
Figure 16: Comparison of error rates $\epsilon$ as a function of bond dimension $\chi$ obtained from the simulation of random nearest-neighbor circuits using MPS and TTN. The results are for a system of $N=32$ qubits with $N_{2g}=310$ two-qubit gates. The inset shows error rates $\epsilon$ as a function of memory footprint $M$ obtained from the same simulation.
@@
Figure 17: Comparison of fidelities $\mathcal$$\tilde${F}} as a function of bond dimension $\chi$ obtained from the simulation of tree-like QAOA circuits using MPS and TTN. The results are for a system of $N=32$ qubits at depth $p=15$ with $N_{2g}=825$ two-qubit gates. The inset shows fidelity $\mathcal$$\tilde${F}} as a function of memory footprint $M$ obtained from the same simulation.
Fig. 15 shows the same comparison, now incorporating a “naive” qubit reordering strategy as Sec. III.2. In both cases, a binary TTN is used, but the strongly correlated qubits are placed together in the TTN. While both approaches benefit from reordering, our DMRG-based algorithm at $\chi=16$ still outperforms Ref. [30] at $\chi=64$. The relative improvement is more pronounced in Ref.[30], indicating that the SVD-based approach is more sensitive to the choice of TTN structure and less robust to long-range entangling gates. This comparison highlights the advantage of our variational framework, which achieves higher accuracy at lower bond dimension while maintaining greater resilience to circuit nonlocality.
@@ Appendix B Comparison between MPS and TTN
We compare MPS and TTN as variational ansatz within the DMRG algorithm for simulating the quantum circuits. All simulations are performed on a system of $N=32$ qubits. For TTN, we adopt a binary tree structure. It is important to note that a direct comparison between TTN and MPS is inherently challenging. As a result, our analysis primarily focuses on the bond dimension $\chi$, which determines the maximum half-chain entanglement entropy that each ansatz can represent [9]. We also provide a comparison of the memory footprint.
@@
Figure 18: Comparison of fidelities $\mathcal$$\tilde${F}} as a function of bond dimension $\chi$ obtained from the simulation of QAOA circuits on random 3-regular graphs using MPS and TTN. The results are for a system of $N=32$ qubits at depth $p=2$ with $N_{2g}=96$ two-qubit gates. The inset shows fidelity $\mathcal$$\tilde${F}} as a function of memory footprint $M$ obtained from the same simulation.